OpenStudy (anonymous):

integral from -1 to 1 of e^(u+1) du.

OpenStudy (anonymous):

\[e^{u+1}=e\times e^u\] is a start

OpenStudy (australopithecus):

use substitution rule s = u +1 now take the derivative of u + 1 ds = du e^(s)ds

OpenStudy (anonymous):

\[e\int_{-1}^1e^udu\] anti derivative is \(e^u\) evaluate at 1 and -1 and subtract

OpenStudy (australopithecus):

then use the fundamental theory F(b) - F(a) where in your case b = 1, and a = -1

OpenStudy (anonymous):

how do you know that e^(u+1)=e * e^u

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

laws of exponents \[b^n\times b^m=b^{n+m}\] \[e^1\times e^u=e^{u+1}\]

OpenStudy (australopithecus):

can't you jsut do a substitution?

OpenStudy (anonymous):

you get as a final answer \(e(e^1-e^{-1})=e^2-1\)

OpenStudy (anonymous):

sure you can, but why bother?

OpenStudy (anonymous):

I am just so confused with this problem, answer in back of book is 0

OpenStudy (australopithecus):

its less complicated

OpenStudy (anonymous):

not zero for sure

OpenStudy (anonymous):

i guess it is a matter of taste you can use a substitution or you can use \(\int cf(x)dx=c\int f(x)dx\) either way

OpenStudy (australopithecus):

ok so use substitution rule s = u +1 now take the derivative of u + 1 ds = du thus we have \[\int\limits_{-1}^{1} e^{s}ds\] the antiderivative of e^(x) = e^(x) so we end up with e^(u+1) as the derivative using the rule F(b) - F(a) we can find the area where a = -1 and b = 1 so e^(1 + 1) - e^(-1 + 1) = e^(2) - 1 I think your book is incorrect https://www.wolframalpha.com/input/?i=ntegral+from+-1+to+1+e^%28u%2B1%29

OpenStudy (anonymous):

okay thanks the substitution makes sense now